OpenStudy (anonymous):

Simplify by rationalizing the denominator 5/square root 7

OpenStudy (anonymous):

\[\sqrt{7}*5/7\]

OpenStudy (tkhunny):

@LS1088 Don't do that. Hints and help. NOT answers. Please read the Code of Conduct. \(\dfrac{5}{\sqrt{7}} = \dfrac{5}{\sqrt{7}}\cdot\dfrac{\sqrt{7}}{\sqrt{7}} = \dfrac{5\sqrt{7}}{\sqrt{7}\cdot\sqrt{7}} = \dfrac{5\sqrt{7}}{7} = \dfrac{5}{7}\sqrt{7}\) Where you stop depends on how the presentation is demanded.

OpenStudy (anonymous):

That isn't the answer

OpenStudy (tkhunny):

Please demonstrate "the answer" and how to achieve it.